OpenStudy (anonymous):

How long would it take $3,500 to grow to $8,900 if the annual rate is 4% and interest in compounded monthly?

OpenStudy (anonymous):

i hate these questions

OpenStudy (anonymous):

me too

OpenStudy (anonymous):

wish i could help haha

OpenStudy (anonymous):

Formula for interest compounded periodically:\[A(t)=P(1+\frac{r}{n})^{nt}\]

OpenStudy (anonymous):

\[8900=3500(1+\frac{.04}{12})^{12t}\implies \frac{89}{35}=(1.0033)^{12t}\]

OpenStudy (anonymous):

\[\implies \frac{ \ln(89)-\ln(35)}{12\ln(1.0033)}=t\]

OpenStudy (anonymous):

Your calculator should be able to give you the number needed for the final answer.

OpenStudy (anonymous):

I think the answer is between 23 and 24 years.

OpenStudy (anonymous):

its 23.6 years